A 6ohm resistance and 8ohm inductance are connected in series then power factor of the circuit is
 

Grade:12

1 Answers

Arun
5 years ago
Dear Khan
 
I think that it can not be in series it is possible in parallel.
 
power factor = 48 / 14 / 10
 
 = 3.42 / 10 = 0.342